x^3/100 = x^log x
Can you solve for ALL values of x, i found that x=10 but can't find the rest of the solutions thank you

Answer by stanbon(75887)

x^3/100 = x^log x
---
x^(3/100) = x^(log(x))
----
log(x) = 3/100
x = 10^(3/100)
x = 100th root of 10^3
x = 100th root of 1000
============================
Cheers,
Stan H.